solve by factoring 8s-11=(s-2)(s-6)
Answer by The One(11)

8s-11=(s-2)(s-6)
8s-11=(s-2)(s-6)
(s-2)(s-6)=8s-11
(s • s+s • -6-2 • s-2 • -6)=8s-11
〖(s〗^2-8s+12)=8s-11
〖(s〗^2-8s+12=8s-11
〖(s〗^2-8s+12-8s)=-11
s^2-16s+12=-11
s^2-16s+12+11=11-11
s^2-16s+23=0
s=(-b±√(b^2-4ac))/2a
where
a=1, b=-16, and c=23
Substitute in the values of a=1, b= -16, and c=23.
s=(-(-16)±√(〖(-16)〗^2-4(1)(23)))/(2(1))
Multiply -1by each term inside the parentheses
s=(16±√(〖(-16)〗^2-4(1)(23)))/(2(1))
Simplify the section inside the radical
s=(16±2√41)/(2(1))
Simplify the denominator of the quadratic formula
s=(16±2√41)/2
First Solve for "+" portion of the "±"
s=(16±2√41)/2
Simplify the expression to solve for the "+" portion of the "±"
s=8+√41
Next, solve the "-" portion of "±"
s=(16-2√41)/2
Simplify the expression to solve for the "-" portion of the "±"
s=8-√41
The final solutions are:
s=8-√41 , s=8+√41
s≈14.40312 or 1.596876